OpenStudy (anonymous):

Solve using matrices. 9x-5y=51 4x + y = 13

hero (hero):

y = 13 - 4x y = (51 - 9x)/-5 y = y -5(13 - 4x) = 51 - 9x -65 +20x = 51 - 9x 29x = 116 x = 4 y = -3